The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 55 May 13, 2001

Contents

Abstracts of 9 refereed papers  1
Abstracts of a non-refereed paper 7
Meeting announcement  7

Abstracts of Refereed Papers

Supernova Remnants in the Southwestern Part of the Small Magellanic Cloud

John R. Dickel^{1}, Rosa M. Williams^{2}, Lynn M. Carter^{1}, D. K. Milne^{3}, Robert Petre^{2},

and Shaun Amy^{3}

1Astronomy Department, University of Illinois at Urbana- Champaign, 1002 West Green Street, Urbana IL 61801, USA 2Goddard Space Flight Center, Code 662, NASA, Greenbelt MD 20771, USA 3 Australia Telescope National Facility, Box 76, Epping NSW 1710, Australia

We have compared radio images from the Australia Telescope Compact Array and x-ray images from ROSAT of an area about 560 pc in diameter around the very extended emission nebula N19 on the southwestern edge of the bright ridge of the Small Magellanic Cloud. These data have allowed us to tentatively identify seven supernova remnants in this region. This number is large enough to suggest that the area has recently undergone some starburst activity. N19 itself appears to contain two overlapping supernova remnants within the extended H II region.

Accepted by: The Astronomical Journal

For preprints, contact johnd@astro.uiuc.edu

Interstellar Bubbles in Two Young H ii Regions

Yaël Nazé^{1,2,5}, You-Hua Chu^{1,6}, Sean D. Points^{1,6}, Charles W. Danforth^{3,6}, Margarita Rosado^{4}, and C.-H. Rosie Chen^{1}

1Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA 2Institut d’Astrophysique et de Géophysique, Avenue de Cointe 5, B 4000 Liège, Belgium 3Department of Physics and Astronomy, Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 4Instituto de Astronomía IA - UNAM, Apartado 70-264, 04510 Mexico D.F., Mexico 5Research Fellow FNRS (Belgium) 6Visiting astronomer, Cerro Tololo Inter-American Observatory

Massive stars are expected to produce wind-blown bubbles in the interstellar medium; however, ring nebulae, suggesting the existence of bubbles, are rarely seen around main-sequence O stars.

To search for wind-blown bubbles around main-sequence O stars, we have obtained high-resolution Hubble Space Telescope WFPC2 images and high-dispersion echelle spectra of two pristine H ii regions, N11B and N180B, in the Large Magellanic Cloud. These H ii regions are ionized by OB associations that still contain O3 stars, suggesting that the H ii regions are young and have not hosted any supernova explosions. Our observations show that wind-blown bubbles in these H ii regions can be detected kinematically but not morphologically because their expansion velocities are comparable to or only slightly higher than the isothermal sound velocity in the H ii regions. Bubbles are detected around concentrations of massive stars, individual O stars, and even an evolved red supergiant (a fossil bubble). Comparisons between the observed bubble dynamics and model predictions show a large discrepancy (1–2 orders of magnitude) between the stellar wind luminosity derived from bubble observations and models and that derived from observations of stellar winds. The number and distribution of bubbles in N11B differ from those in N180B, which can be explained by the difference in the richness of stellar content between these two H ii regions. Most of the bubbles observed in N11B and N180B show a blister-structure, indicating that the stars were formed on the surfaces of dense clouds. Numerous small dust clouds, similar to Bok globules or elephant trunks, are detected in these H ii regions and at least one of them hosts on-going star formation.

Accepted by: The Astronomical Journal

For preprints, contact naze@astro.uiuc.edu or naze@astro.ulg.ac.be
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0104420

HST/NICMOS detection of a partially embedded, intermediate-mass pre-main-sequence population in the 30 Doradus Nebula

Wolfgang Brandner^{1,2}, Eva K. Grebel^{3}, Rodolfo H. Barbá^{4}, Nolan R. Walborn^{5}, Andrea Moneti^{6}

1 European Southern Observatory, Karl-Schwarzschild-Str. 2, D-85748 Garching, Germany 2 University of Hawaii, Institute for Astronomy, 2680 Woodlawn Dr., Honolulu, HI 96822, USA 3 Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 4 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de la Plata, Paseo del Bosque S/N, 1900 La Plata, Argentina 5 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 6 Institut d’Astrophysique Paris, 98bis Blvd Arago, F-75014 Paris, France

We present the detection of an intermediate-mass pre-main-sequence population embedded in the nebular filaments surrounding the 30 Doradus region in the Large Magellanic Cloud (LMC) using HST/NICMOS. In addition to four previously known luminous Class I infrared “protostars,” the NICMOS data reveal 20 new sources with intrinsic infrared excess similar to Galactic pre-main sequence stars. Based on their infrared brightness, these objects can be identified as the LMC equivalent of Galactic pre-main sequence stars. The faintest LMC Young Stellar Objects in the sample have colors similar to T Tauri and have about the same brightness as T Tauri if placed at the distance of the LMC. We find no evidence for a lower-mass cut-off in the initial mass function. Instead, the whole spectrum of stellar masses from pre-main sequence stars with ∼1.5 M_{⊙} to massive O stars still embedded in dense knots appears to be present in the nebular filaments. The majority of the young stellar objects can be found to the north of the central starburst cluster R136. This region is very likely evolving into an OB association.

Accepted by: The Astronomical Journal (August 2001 issue) For preprints, contact wbrandne@eso.org

Also available from the URL http://arXiv.org/abs/astro-ph/0104274

Stellar evolution with rotation VII: Low metallicity models and the blue to red supergiant ratio in the SMC

A. Maeder and G. Meynet

Geneva Observatory CH–1290 Sauverny, Switzerland

We calculate a grid of models with and without the effects of axial rotation for massive stars in the range of 9 to 60 M_{⊙} and metallicity Z = 0.004 appropriate for the SMC. Remarkably, the ratios Ω/Ω_{crit} of the angular velocity to the break–up angular velocity grow strongly during the evolution of high mass stars, contrary to the situation at Z = 0.020. The reason is that at low Z, mass loss is smaller and the removal of angular momentum during evolution much weaker, also there is an efficient outward transport of angular momentum by meridional circulation. Thus, a much larger fraction of the stars at lower Z reach break–up velocities and rotation may thus be a dominant effect at low Z. The models with rotation well account for the long standing problem of the large numbers of red supergiants observed in low Z galaxies, while current models with mass loss were predicting no red supergiants. We discuss in detail the physical effects of rotation which favour a redwards evolution in the HR diagram. The models also predict large N enrichments during the evolution of high mass stars. The predicted relative N–enrichments are larger at Z lower than solar and this is in very good agreement with the observations for A–type supergiants in the SMC.

Accepted by: Astronomy & Astrophysics

For preprints, contact Andre.Maeder@obs.unige.ch
Also available from the URL http://xxx.sissa.it/abs/astro-ph/0105051

56Ni Mixing in the Outer Layers of SN 1987A

Robert C. Mitchell^{1}, E. Baron^{1}, David Branch^{1}, Peter Lundqvist^{2}, Sergei Blinnikov^{3},

Peter H. Hauschildt^{4}, and Chun S. J. Pun^{5}

1 Physics and Astronomy Dept, University of Oklahoma, 440 West Brooks, Norman, OK 73019-0261, USA 2 Stockholm Observatory, S–133 36 Saltsjöbaden, Sweden 3 Institute of Theoretical and Experimental Physics, 117218, Moscow, Russia 4 Dept of Physics and Astronomy & Center for Simulational Physics, University of Georgia, Athens, GA 30602, USA 5 Laboratory for Astronomy and Solar Physics, NASA/GSFC, Greenbelt, MD 20771, USA

Supernova 1987A remains the most well-observed and well-studied supernova to date. Observations produced excellent broad-band photometric and spectroscopic coverage over a wide wavelength range at all epochs. Here, we focus on the very early spectroscopic observations. Only recently have numerical models been of sufficient detail to accurately explain the observed spectra. In SN 1987A, good agreement has been found between observed and synthetic spectra for day one, but by day four, the predicted Balmer lines become much weaker than the observed lines. We present the results of work based on a radiation-hydrodynamic model by Blinnikov and collaborators. Synthetic non-LTE spectra generated from this model by the general radiation transfer code PHOENIX strongly support the theory that significant mixing of ^{56}Ni into the outer envelope is required to maintain strong Balmer lines. Preliminary results suggest a lower limit to the average nickel mass of 1.0 x 10^{−5} solar masses is required above 5000 km s^{−1} by day four. PHOENIX models thus have the potential to be a sensitive probe for nickel mixing in the outer layers of a supernova.

Accepted by: The Astrophysical Journal

For preprints, contact mitchell@mail.nhn.ou.edu
Also available from the URL xxx.lanl.gov/abs/astro-ph/0104148

A 168-s X-Ray Pulsar in the SMC Observed with ASCA

Jun Yokogawa^{1}, Ken’ichi Torii^{2}, Takayoshi Kohmura^{3}, and Katsuji Koyama^{1}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto 606-8502, Japan 2 RIKEN (Institute of Physical and Chemical Research), 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 3 Department of Earth and Space Science, Graduate School of Science, Osaka University, 1-1 Machikaneyama-cho, Toyonaka, Osaka, 560-0043, Japan

We report on the detection of 167.8 ± 0.2 s pulsations from AX J0052.9−7157 in the Small Magellanic Cloud with ASCA. The pulsations show an energy-independent sinusoidal profile. The X-ray spectrum is described by an absorbed hard power-law (photon index ∼ 0.7) with a luminosity of ∼ 8.8 × 10^{35} erg s^{−1} at a distance of 60 kpc. The position is determined to be (R.A., Dec.)_{J2000} = (00^{h}52^{m}59^{s}.2, −71^{∘}57′58′′), with an error radius of ∼ 30′′. We have found that a variable Be/X-ray binary RX J0052.9−7158 is located within the 30′′ error circle, and is also located at the edge of the 10′ error circle of the 169-s pulsar XTE J0054−720. We therefore conclude that these are the same source, a Be/X-ray binary pulsar.

Accepted by: PASJ

For preprints, contact jun@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www-cr.scphys.kyoto-u.ac.jp/member/jun/job/

A Method to Determine the Star Formation History of a Mixed Stellar Population

Jason Harris^{1} and Dennis Zaritsky^{2}

1 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 2 Steward Observatory, 933 E. Cherry Ave., Tucson, AZ 85721, USA

We present a method to determine the star formation history of a mixed stellar population from its photometry. We perform a chi-squared minimization between the observed photometric distribution and a model photometric distribution, based on theoretical isochrones. The initial mass function, distance modulus, interstellar reddening, binary fraction and photometric errors are incorporated into the model, making it directly comparable to the data. The model is a linear combination of individual synthetic color-magnitude diagrams (CMDs), each of which represents the predicted photometric distribution of a stellar population of a given age and metallicity. We present extensive tests of the algorithm, using both simulated and actual photometric data. From a preliminary application of the algorithm to a subregion of the Large Magellanic Cloud (LMC), we find that the that the lull in star formation observed among the LMC’s cluster population between 3 and 8 Gyr ago is also present in the field population. The method was designed with flexibility and generality in mind, and we make the code available for use by the astronomical community.

Accepted by: Astrophysical Journal Supplement Series For preprints, contact jharris@stsci.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0104202

Cepheid Mass-Luminosity Relations from the Magellanic Clouds

J. P. Beaulieu^{1}, J.R. Buchler^{2}, Z. Kollath^{3}

1 Institut d’Astrophysique de Paris, Paris, France 2 University of Florida, Gainesville, USA ^{3} konkoly Observatory, Budapest, Hungary

The OGLE data base is used in conjunction with Kurucz atmosphere models to generate sets of period, effective temperature and luminosity for fundamental and overtone Magellanic Cloud Cepheids. The Florida pulsation code (with linear turbulent convection) is then used to compute masses for these stars, assuming an average composition of (X=0.716, Z=0.010) for the LMC and of (X=0.726, Z=0.004) for the SMC. The average M–L relation for the fundamental Cepheids matches closely that for the first overtone Cepheids for each Magellanic Cloud. Neither the SMC nor the LMC average log M–log L relations are straight, but have a noticeable curvature. In view of the uncertainties in distance and reddening we have adopted three different choices for these quantities. The results based on the ‘long’ distance scale to the clouds give a better agreement between theory and and observations than the ‘short’ one. All the current evolutionary tracks predict systematically larger masses for given luminosities than our observationally derived ones, especially at the high end. Moreover, our study confirms that the evolutionary tracks of the low mass stars in SMC are not in agreement with the observations as they do not extend sufficiently blueward and do not penetrate deep enough into the instability strip, or not at all. The inference of masses directly from the observational database yields a novel and strong constraint on evolutionary calculations.

Accepted by: Astronomy and Astrophysics

For preprints, contact beaulieu@iap.fr, buchler@phys.ufl.edu, kollath@konkoly.hu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0010417

The Distance to the Magellanic Clouds

Bohdan Paczynski^{1}

1 Princeton University Observatory, Peyton Hall, NJ 08544, USA

The distance to LMC and SMC is a subject of controversy, with the difference between the extreme values in distance moduli exceeding 0.5 mag. While currently the best calibrated method is based on red clump giants, and the near future improvement is most likely to come from detached eclipsing binaries, the ultimate goal is to have a purely geometrical determination. The best prospect will be to use relatively wide binary stars, for which spectroscopic orbits will be obtained with large ground based telescopes, and astrometric orbits will be obtained either with SIM, or with future ground based interferometers. A preliminary list of 25 candidate systems is presented. It is based on OGLE catalogs of BVI photometry.

Submitted to: Acta Astronomica

For preprints, contact bp@astro.princeton.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0104182

Abstracts of Non-Refereed Papers

Superbubble metallicities and X-ray luminosities

Sergiy Silich^{1,2} and Sally Oey^{3}

1 Instituto Nacional de Astrofísica Optica y Electrónica, AP 51 y 216, 72000 Puebla, México 2 Main Astronomical Observatory National Academy of Sciences of Ukraine, 03680, Kyiv-127, Golosiiv, Ukraine 3 Lowell Observatory, 1400 West Mars Hill Rd., Flagstaff, AZ 86001, USA

The LMC harbors a wealth of superbubbles, many of which have been detected in soft X-rays. The comparison of the observed X-ray luminosities with the model predictions revealed that some X-ray bright superbubbles have X-ray luminosities up to an order of magnitude higher than the expected values. Two models have been proposed to resolve this discrepancy: off-center SN explosions (CM90) and mass-loading by dense cloud destruction.

Here we examine another possibility, namely time variation in the metallicity of the X-ray emitting gas, that results from metal ejection by SN explosions of the parent OB-association. As an example, we use the LMC superbubble DEM L50.

Our results imply that, if the host galaxy metal abundance is well below solar (Z_{ism} < 0.1Z_{⊙}), then enrichment by massive star products may completely dominate the interior bubble chemical composition soon after the onset of SN explosions. This enhances superbubble X-ray luminosities and predicts only weak dependence on the parent galaxy metallicity.

To appear in: Extragalactic Star Clusters, IAU Symposium 207, eds. E.K. Grebel, D. Geisler, D. Minniti (Provo: ASP) For preprints, contact silich@inaoep.mx

Meeting Announcement

Galaxies: The Third Dimension

Instituto de Astronomía, Universidad Nacional Autónoma de México December 3–7, 2001

Scientific goals and topics

This meeting is intended to bring together researchers who use or develop 3D spectroscopic instruments (such as integral field spectrographs and Fabry-Perots). We will exchange ideas on the application of 3D methods to the study of extragalactic astronomy. A particular emphasis will be given to presentations concerning new instruments especially for use on large diameter telescopes. A particular focus will involve applications in extragalactic research. The meeting will include approximately 10 review talks and a number of shorter talks on the different topics of the scientific program.

1. 3D Instruments & techniques for large diameter telescopes.

2. Interrelationship between gas and stars in galaxies. (star formation, bubbles, superbubbles and supershells, shocks).

3. Dynamics of isolated galaxies.

4. Dynamics of interacting galaxies: pairs, compact groups and clusters.

5. Study of the central regions of AGNs.

Scientific Organizing Committee

G. Monnet, R. Bacon, T.A. Lozinskaya, J. Sulentic, C. Esteban, Y.H. Chu, J.J. Gonzalez, J. Boulesteix, B. Tully, E. Brinks, L. Aguilar, S. Silich, S. Veilleux, P. Palunas, J. Franco, J. Guichard, Margarita Rosado (Chair).

Local Organizing Committee

P. Ambrocio-Cruz, L. Arias, A. Bullejos, E. De la Fuente, I. Fuentes, I. Puerari, H. Plana, M. Valdez-Gutierrez, A. Bernal, L. Binette, M. Rosado.

Registration

You are invited to pre-register NOW for this event by using our on-line registration program on our

website at http://www.astroscu.unam.mx/3Dgal/.

Further announcements will be sent to all those who pre-register. Report any problem when registering to the registration manager at 3Dgal@astroscu.unam.mx.

The second announcement, detailing the definitive location of the meeting, the registration fee, the list of invited speakers, the publisher company, etc., will be sent before the end of June. Please pre-register now to become part of our mailing list. All the information contained in this e-mail, and future announcements, will remain available at our website above.

Contact Address

Dedicated Fax: (US) 1-209-396 8367 (no cover page needed)

http://www.astroscu.unam.mx/3Dgal/

Postal address:

Margarita Rosado

Galaxies: The Third Dimension

Instituto de Astronomía, UNAM

Apartado Postal 70-264

04510 México, D.F.

México